A 50 g of Cupric Sulfate has been added to a 300g of 2% Cupric Sulfate solution at 10 oC.
How many Cupric Sulfate will not be dissolved? What kind of solution will be formed? If
unsaturated, how many more can be dissolved?
1
Expert's answer
2020-10-21T13:17:46-0400

Solubility of CuSO4 at 10 "C^\circ" is 17.2 g/100 g water

mass percent of CuSO4 in such solition is

"w(CuSO_4)= \frac{17.2}{17.2+100}=0.147"


1) Find mass of CuSO4 in 300 g 2% CuSO4 solution

"m(CuSO_4) = 300\times0.02=6 g"


2) Chech up how many grams of CuSO4 one needs to add to the 300 g of 2% solution for it to become saturated:

let x to be the mass of the CuSO4 which was added

"0.147 = \frac{6+x}{300+x}"

"x= 44.67 g"

So, the solution will be saturated if we add 44.67 g of CuSO4.

If we add 50 grams of CuSO4, the solution also will be saturated and some CUSO4 will not be dissolved:

"50-44.67=5.33 g"